# ViTextVQA: A Large-Scale Visual Question Answering Dataset for Evaluating Vietnamese Text Comprehension in Images

### v2

@@ Abstract
Visual Question Answering (VQA) is a complicated task that requires the capability of simultaneously processing natural language and images. Initially, this task was researched, focusing on methods to help machines understand objects and scene contexts in images. However, some text appearing in the image that carries explicit information about the full content of the image is not mentioned. Along with the continuous development of the AI era, there have been many studies on the reading comprehension ability of VQA models in the world. As a developing country, conditions are still limited, and this task is still open in Vietnam. Therefore, we introduce the first large-scale dataset in Vietnamese specializing in the ability to understand text appearing in images, we call it ViTextVQA (Vietnamese Text-based Visual Question Answering dataset) which contains over 16,000 images and over 50,000 questions with answers. Through meticulous experiments with various state-of-the-art models, we uncover the significance of the order in which tokens in OCR text are processed and selected to formulate answers. This finding helped us significantly improve the performance of the baseline models on the ViTextVQA dataset. Our dataset is available at this link for research purposes.
@@ 1 Introduction
In recent years, Visual Question Answering (VQA) has received special attention from researchers in the fields of Computer Vision (CV) and Natural Language Processing (NLP). The appeal of VQA is boosted when Flamingo [1], GPT4 [2], Gemini [3] - the most powerful chatbots currently integrated with question and answer features on images. Therefore, this field is on a strong development path around the world, especially in rich resource languages such as English and Chinese. Many VQA datasets such as [4, 5, 6, 7] have been released, and new methods with outstanding performance continuously developed, making this multimodal learning with remarkable advancements. Neural and transformer architectures play a pivotal role in a diverse array of tasks that extend beyond traditional methods such as [8, 9, 10].
This task requires a flexible combination of image processing and language processing to address the challenges associated with answering questions based on visual information. In VQA, the model needs to understand not only the content of an image, video, or document image but also the meaning of the question. The model must then generate a meaningful and relevant answer to that question. This places high demands on a deep understanding of both natural language and image analysis capabilities and requires the model to be able to synthesize diverse information from different information sources.
In the context of Vietnamese VQA, Tran et al. [11] built the ViVQA dataset, the first dataset for research on the VQA task for Vietnamese. The ViVQA dataset with 15,000 samples is a small part of the VQA v2 dataset as it was constructed using the semi-automatic method. However, the publication of this dataset is not convincing enough. The shortcomings of the dataset have been elaborately presented by the research team of Nguyen et al. [12]. Therefore, in 2023, they launched a dataset for the open-ended VQA task in Vietnamese called OpenViVQA. This also opens up a new direction of research as well as a different form of the VQA task, which has open-ended questions and open-ended answers in Vietnamese. The dataset includes 11,199 images together with 37,914 question-answer pairs (QAs) created manually.
While the ViVQA dataset does not mention scene text, the OpenViVQA dataset has half of the total samples that exploit this element as one of its challenges. However, OpenViVQA also brings with it some limitations, creating an opportunity for the development of a new dataset that focuses specifically on this aspect.
There are several reasons why OpenViVQA has limitations. First, the number of samples in the dataset may not be large enough to thoroughly and reliably evaluate the performance of the model in scene text. Second, the diversity of scene text in the dataset may be limited, unable to fully reflect the real-life situations the model may encounter. Finally, evaluation capabilities on the OpenViVQA dataset may not reach the level of detail and rigor for scene text processing capabilities when the answers to questions in this dataset are in free form.
@@
To address the above limitations, we developed a new dataset called Vietnamese Text-based Visual Question Answering (ViTextVQA), which includes 16,762 images and 50,342 question-answer pairs. This dataset focuses mainly on capturing information from text and scene text appearing in images to address the limitations that traditional VQA models often encounter. ViTextVQA is not only a diverse dataset but also creates a basis for evaluating and improving VQA models capable of handling text, especially in the context of the Vietnamese language.
Our main contributions include the following:
  • Create the first high-quality large-scale dataset for text-based VQA task in Vietnamese, focusing on scene text and text appearing in the image.
  • Analyze the challenge of the ViTextVQA dataset by evaluating the performance of the OCR system.
  • Through our extensive experiments, we found that VQA models using ViT5 as their backbone behave as the answer selector methods when OCR text is suffixed for the question.
  • Demonstrate the important role of OCR text on the performance of VQA models on the ViTextVQA dataset, especially arrangements. Our findings underscore the effectiveness of the top-left to the bottom-right sort, resulting in remarkable enhancements in the performance.
This article is organized explicitly as follows: Section 2 presents an overview of studies related to our research field. In Section 3, we introduce the data creation process and some detailed analysis of our dataset. Next, Section 4 introduces some baselines that we conduct experiments with. Section 5 details the design of the experiments and the evaluation of the results, including comparisons with the baselines. Section 6 analyzes in detail the results achieved. Section 7 discusses some Vietnamese characteristics. Finally, Section 8 summarizes the research and suggests future directions.
@@ 2.1.1 Former Visual Question Answering Datasets
Dataset Type Image Source Annotation Method Language DAQUAR Non text NYU-DepthV2 Human annotation English VQA v1 Non text MS COCO Human annotation English VQA v2 Non text MS COCO Human annotation English TextVQA Text MS COCO Human annotation English OCR-VQA Text Book covers, movie covers Semi-Automation English ST-VQA Text Multiple sources Human annotation English DocVQA Text Scanned documents, receipts etc. Human annotation English VisualMRC Text Internet sources Human annotation English
There are a lot of large-scale and high-quality VQA datasets in the world, especially in the English language. These datasets are not only valuable resources but also crucial sources of inspiration to assist us in constructing the ViTextVQA dataset. Based on research and VQA datasets in English, we can learn and apply effective methods and techniques in the process of developing our Vietnamese VQA dataset. More details of these datasets are in Table 1.
DAQUAR dataset: Malinowski and Fritz [13] introduced the DAtaset for QUestion Answering on Real-world images (DAQUAR) dataset as an early VQA benchmark, initiated the Visual Turing Challenge. They introduced this dataset as a benchmark for evaluating the performance of VQA models. It aimed to assess the capability of models to process both image data and natural language questions, showcasing their ability to bridge the gap between vision and language understanding. The dataset includes a variety of question types, ranging from simple factual queries to more complex questions that necessitate deeper understanding and reasoning about the visual content. This diversity aims to challenge models across different aspects of visual and linguistic comprehension.
VQA v1 dataset: Antol et al. [14] first proposed the free-form question in the VQA task and published the Visual Question Answering v1 (VQA v1) dataset for this task. This dataset has opened a major revolution in natural language processing and has attracted a large number of researchers. During those years, there have been many methods proposed to solve the problem, typically Multimodal Compact Bilinear Pooling (MCB) [15], Multimodal Residual Learning (MUTAN) [16], Neural Module Network (NMN) [17].
VQA v2 dataset: The VQA v1 dataset still has shortcomings that researchers need to pay attention to. The question types were repeated and the answers were only based on simple image features, which led to VQA models at that time only focusing on learning image features and not understanding question content and visual information. To improve visibility for training models, Goyal et al. [18] published the Visual Question Answering v2 (VQA v2) dataset by adding images to the original dataset of Antol et al. [14] and the questions link not just to a single image but to a pair of similar images that lead to two different answers to that question.
TextVQA dataset: Traditional VQA task often ignore text content appearing in images, or they are not capable of reading text appearing in images. To expand the scope of this task, Singh et al. [4] built the TextVQA dataset, which contains 45,336 questions on 28,408 images, requiring questions to use text appearing in the image as the answer. At the same time, they also presented the LoRRA model, which added OCR to retrieve text appearing in images to use as answers.
OCR-VQA dataset: Realize that VQA methods often do not consider text processing in images, which can reduce system performance. To solve this problem, Mishra et al. [19] proposed a new task called Optical Character Recognition - Visual Question Answering (OCR-VQA). The goal of this task is to answer questions about the image by reading and understanding the text it contains. To support research on this task, the article introduces a large dataset called OCR-VQA-200K, including 207,572 book cover images and more than 1 million corresponding question-answer pairs using the semi-automation annotation method.
ST-VQA dataset: VQA datasets and models at the time had serious limitations in ignoring the content of the text in the image. Therefore, when the questions required to understand scene text, VQA models led to disappointing results. To address a significant drawback in VQA datasets and models, [5] introduced the Scene Text VQA (ST-VQA) dataset, specifically designed to tackle the challenge of understanding and answering questions related to texts that appear in the image. The dataset includes question-answer pairs formulated in a way that requires knowledge of the text present in the image for accurate answers. This initiative aimed to improve the performance of models in scenarios where scene text comprehension is crucial.
DocVQA dataset: Unlike existing VQA datasets that focus on natural scene images, DocVQA introduced by Mathew et al. [20] uses images of real-world documents like invoices, receipts, contracts, etc. This presents unique challenges, as the information is often presented in structured formats like tables, forms, and lists, requiring models to reason beyond simple object recognition. DocVQA comprises 50,000 question-answer pairs based on 12,767 document images. These images come from various industry documents and prioritize pages with tables, forms, lists, and figures. This dataset also serves as a benchmark for researchers to develop and evaluate VQA models that can handle document images effectively. Overall, DocVQA provides a valuable resource for advancing research in VQA and its applications in document understanding and information retrieval.
VisualMRC dataset: Tanaka et al. [21] presented a significant advancement in the field of VQA on document images, surpassing the scope of previous datasets like DocVQA by introducing an abstractive novel task. Unlike traditional question-answering datasets, VisualMRC focuses on machine reading comprehension, challenging models to not only extract information from document images but also comprehend and reason about the content to generate abstractive answers. With a substantial dataset comprising over 30,000 question-answer pairs and over 10,000 diverse document images across various web pages, VisualMRC provides a rich and expansive resource for training and evaluating models. The dataset includes open-ended questions, requiring models to generate natural language responses based on a deep understanding of the document layout, text, and visual elements.
@@ 2.1.2 Visual Question Answering Datasets in Vietnamese
Dataset Type Image Source Annotation Method Language ViVQA Non text MS COCO Semi-Automation Vietnamese EJVVQA Non text Internet sources Human annotation English, Japanese, Vietnamese OpenViVQA Text & Non text Internet sources Human annotation Vietnamese ViTextVQA(ours) Text Internet sources, hand-craft Human annotation Vietnamese
We not only refer to English sources but also to current available VQA Vietnamese datasets when developing ViTextVQA. Referencing several Vietnamese datasets, we better understand the language structure and specific content of Vietnamese, helping to create questions and answers that accurately reflect the cultural context and characteristics of Vietnam. This not only supplements and improves these Vietnamese datasets but also helps overcome their limitations. More details of these datasets are in Table 2.
ViVQA dataset: Tran et al. [11] built the first VQA dataset for the task of automatically asking and answering images based on Vietnamese called Vietnamese Visual Question Answering (ViVQA) dataset. With 10,320 images and 15,000 question-answer pairs, ViVQA meets a large enough number to investigate and develop methods to solve VQA task in Vietnamese. Because the dataset is created using a semi-automatic method, the questions do not really have the nature of the Vietnamese language, and there are no text elements in the image that are mentioned.
EVJVQA dataset: The EVJVQA dataset, introduced in the work by Nguyen et al. [22], stands as a unique compilation of multilingual visual question-answer pairs meticulously curated through human annotation, utilizing images captured in the picturesque landscapes of Vietnam. Encompassing a total of approximately 33,790 question-and-answer pairs alongside a repository of over 5,000 associated images, EVJVQA offers a multifaceted and immersive reservoir of information. Notably, the dataset comprises 11,689 question-answer pairs in Vietnamese, with the remaining entries expertly translated into English and Japanese from this original Vietnamese subset. This expansive dataset has found practical application as the primary data source for the Multilingual Visual Question Answering task in the 9th workshop on Vietnamese Language and Speech Processing (VLSP 2022111https://vlsp.org.vn/vlsp2022/eval/evjvqa), showcasing its significance in advancing research and development in the field.
OpenViVQA dataset: The existing VQA dataset in the Vietnamese language constructed by Tran et al. [11] is addressed as an answer selection or answer classification task. Realize the limitations of ViVQA dataset, Nguyen et al. [12] created the OpenViVQA dataset, the first large-scale dataset for VQA with open-ended answers in Vietnamese, consisting of 11,199 images associated with 37,914 question-answer pairs, In this dataset, answers are created in open-ended answer format, which means the answer is independent of any rule, they are generated in a variety of ways using natural language forms such as words, sentences or phrases. Therefore, this dataset can not be used to conduct experiments as a classification task.
As can be seen in the Vietnamese VQA datasets above reveals certain limitations. Particularly, the ViVQA dataset appears to be merely a translation from English. Furthermore, the EVJVQA dataset does not leverage scene text, thus missing out on potentially valuable information. Evaluating the efficacy of this aspect in the OpenViVQA dataset is challenging due to the nature of its free-form answer. These limitations underscore the further development opportunities, highlighting the potential for improvement in our ViTextVQA dataset.
@@ 2.2.3 Vision Transformer Model - Transformer Language Model Based Method
3 ViTextVQA Dataset
We conducted a rigorous process of data annotation take over 6 weeks, involving the participation of sixteen native Vietnamese speakers. The process of building the dataset is presented in Section 3.1. The prominent and intriguing features of the dataset, which we have identified through the labeling process, are detailed in Section 3.2. This Section includes an in-depth analysis of elements such as Question, Answer, and Object.
@@ 3.1.3 Question-Answer Creation
Next, we recruited sixteen undergraduate students. Then, we trained them to understand the annotation guideline and asked them to annotate on about 100 test images. After ensuring that all students understood and applied the annotation guideline correctly, we appointed one person from our team for every four students to act as a supervisor. Annotation supervisors are responsible for monitoring and supporting the annotators, ensuring that everyone is performing according to the required process and quality.
@@ 3.2 Dataset Analysis
To further analyze the characteristics of the ViTextVQA dataset, we compiled statistics on the distribution of the lengths of questions and answers. In addition to that, we also gathered information on Part-of-Speech (POS) and Named Entity Recognition (NER). POS analysis reveals the grammatical structure and meaning of both questions and answers, while NER provides an overview of how words are used in questions and answers. Therefore, we utilized VnCoreNLP [59] - a powerful tool for processing Vietnamese text - to extract POS tags and NER tags. Our statistical overview is presented in Section 3.2.1 and Section 3.2.2.
3.2.1 Question Analysis
Question Length
@@ 3.2.1 Question Analysis
POS Tags in Questions
In VQA tasks, analyzing the Part-of-Speech (POS) tags of words in questions is crucial for understanding their grammatical structure and meaning, providing insights into their overall grammatical makeup.
Figure 7: Distribution of Postag in question.
Figure 8: Distribution of NER in question.
As can be seen in Figure 8, our dataset is annotated using various types of POS tags in question. Among them, Noun (N) appears the most, with more than 140,000 cases, usually referring to objects or people in the image. The Pronoun (P) is used to replace the noun and appears over 61,000 times, playing an important role in expressing the relationship between the components in the question.
Verb (V) is another important part of questions, with approximately 75,000 cases, often describing the action or state of the object in the image. Additionally, the Preposition (E) appears less than half that of the Verb (V) approximately 35,000 times, commonly used to determine relationships between location, time, etc. between the components in the question.
Adjectives (A) and Classifier noun (Nc) appear relatively frequently, with both types of words appearing more than 13,000 times. In the Vietnamese language, Nc represents nouns with classifier properties, while Adjectives (A) are used as a description of their characteristics. The Adverb (R) appears just more than 6,100 times, often used to add information about the manner, extent, time, or frequency of an action, event, or adjective to make the question more vivid.
Other types of words such as Numeral/Quantity (M), Determiner (L), Subordinating conjunction (C), etc. also play an important role in identifying and describing details of images. Although they only make up a small part of the text, they help annotators create diverse grammatical structures and create challenges for VQA models.
NER Tags in Question
Like POS tags, Named Entity Recognition (NER) also provides an overview of language usage in questions, with special attention to entities such as places, organizations, personal entities, and other entities.
Figure 8 shows that the majority of the words in the question of the dataset are not classified as specific entities with more than 350,000 “O” tags which may be common words related to objects such as “person”, “object”, or “action”. Additionally, specific entities such as locations (B-LOC), organizations (B-ORG, I-ORG), individuals (B-PER), and other entities (B-MISC, I-MISC) account for a very small part. This difference can be explained by the fact that questions are not too long and often ask about only part of the text appearing in the image.
3.2.2 Answer Analysis
Answer Length

### v3

# ViTextVQA: A Large-Scale Visual Question Answering Dataset and an Novel Multimodal Feature Fusion Method for Vietnamese Text Comprehension in Images

@@ Abstract
Visual Question Answerinng (VQA) is a complicated task that requires the capability of simultaneously processing natural language and images. This task was initially researched with a focus on developing methods to help machines understand objects and scene contexts in images. However, some scene text that carries explicit information about the full content of the image is not mentioned. Along with the continuous development of the AI era, there have been many studies on the reading comprehension ability of VQA models in the world. Therefore, we introduce the first large-scale dataset in Vietnamese specializing in the ability to understand scene text, we call it ViTextVQA (Vietnamese Text-based Visual Question Answering dataset) which contains over 16,000 images and over 50,000 questions with answers. To tackle this task efficiently, we propose ViTextBLIP-2, an novel multimodal feature fusion Method, which optimizes Vietnamese OCR-based VQA by integrating a frozen Vision Transformer, SwinTextSpotter OCR, and ViT5 LLM with a trainable Q-Former for multimodal feature fusion. Through experiments with various state-of-the-art models, we uncover the significance of the order in which tokens in OCR text are processed and selected to formulate answers. This finding helped us significantly improve the performance of the baseline models on the ViTextVQA dataset. Our dataset is available  [^1] https://github.com/minhquan6203/ViTextVQA-Dataset for research purposes.
@@ 1 Introduction
In recent years, Visual Question Answering (VQA) has garnered significant attention from researchers in Computer Vision (CV) and Natural Language Processing (NLP). The appeal of VQA has surged with the advent of powerful models like Flamingo [1], GPT-4 [2], and Gemini [3], which integrate advanced image-based question-answering capabilities. This has propelled VQA into a phase of robust global development, particularly in resource-rich languages such as English and Chinese. Numerous VQA datasets, including [4, 5, 6, 7], have been released, accompanied by innovative methods showcasing remarkable performance, driven largely by neural and transformer architectures [8, 9, 10], which extend beyond traditional approaches.
In the realm of Vietnamese VQA research, Tran et al. [11] introduced ViVQA, the first dataset tailored for this task, comprising 15,000 samples derived semi-automatically from VQA v2. Despite its pioneering role, ViVQA faced scrutiny for its limitations, as later detailed by Nguyen et al. [12]. Addressing these issues, the OpenViVQA dataset emerged in 2023, designed for open-ended Vietnamese VQA with 11,199 images and 37,914 manually crafted question-answer (QA) pairs. This dataset marked a shift toward open-ended questions and answers, opening new research avenues. Subsequently, the ViOCRVQA dataset [13] scaled up efforts with nearly 30,000 images and over 120,000 QA pairs, emphasizing text-based question answering in images.
While ViVQA largely overlooks scene text, OpenViVQA incorporates it in about half its samples, and ViOCRVQA fully focuses on text-based VQA. However, these datasets still exhibit limitations. First, their sample sizes may not suffice for a comprehensive evaluation of scene text understanding. Second, the diversity of text in images might not fully capture real-world complexity. Lastly, OpenViVQA’s free-form answers complicate objective performance assessment, while ViOCRVQA, despite its scale, reveals gaps in addressing specialized needs.
@@
To overcome these shortcomings, we introduce the Vietnamese Text-based Visual Question Answering (ViTextVQA) dataset, featuring 16,762 images and 50,342 QA pairs. ViTextVQA prioritizes extracting information from scene text to tackle challenges traditional VQA models face (see Figure 1). Beyond offering diversity, it lays a foundation for evaluating and enhancing VQA models adept at handling text, particularly in Vietnamese. To further advance this task, we propose ViTextBLIP-2, an extension of BLIP-2, optimized for Vietnamese text-based VQA by integrating pretrained vision and language models with a trainable Q-Former, as detailed in Section 4.
Our main contributions are described as follows:
  • First, we create the first high-quality, large-scale dataset for text-based VQA in Vietnamese, emphasizing scene text in images.
  • Then, we conduct analysis of ViTextVQA challenges through OCR system performance evaluation, underscoring its pivotal role in VQA models. Our findings reveal that sorting text from top-left to bottom-right markedly boosts performance.
  • Finally, we propose ViTextBLIP-2, a novel method enhancing Vietnamese VQA efficiency, leveraging frozen pretrained components and a trainable Q-Former for multimodal integration.
This article is structured as follows: Section 2 reviews related studies. Section 3 details the ViTextVQA creation process and analysis. Section 4 our proposed ViTextBLIP-2. Section 5 presents baseline models and outlines experimental design and results. Section 6 delves into result analysis. Section 7 explores Vietnamese-specific traits. Finally, Section 8 concludes with future directions.
@@ 2.1.1 Former Visual Question Answering Datasets
Dataset Type Image Source Annotation Method Language DAQUAR [14] Non text NYU-DepthV2 Human annotation English VQA v1 [15] Non text MS COCO Human annotation English VQA v2 [16] Non text MS COCO Human annotation English TextVQA [4] Text MS COCO Human annotation English OCR-VQA [17] Text Book covers, movie covers Semi-Automation English ST-VQA [5] Text Multiple sources Human annotation English DocVQA [18] Text Scanned documents, receipts etc. Human annotation English VisualMRC [19] Text Internet sources Human annotation English
The field of Visual Question Answering (VQA) has been significantly advanced by several large-scale, high-quality datasets in English, which serve as valuable resources and inspiration for our ViTextVQA dataset development. The DAQUAR dataset, introduced by Malinowski and Fritz [14], was an early VQA benchmark that initiated the Visual Turing Challenge, assessing models’ abilities to process both image data and natural language questions. Antol et al. [15] revolutionized the field by proposing free-form questions in the VQA v1 dataset, attracting numerous researchers and methodologies such as MCB, MUTAN, and NMN. To address shortcomings in VQA v1, Goyal et al. [16] published VQA v2, enhancing the dataset with paired similar images leading to different answers for the same question. Recognizing the importance of scene text, Singh et al. [4] developed the TextVQA dataset with 45,336 questions requiring scene text as answers, while also introducing the LoRRA model incorporating OCR capabilities. Similarly, Mishra et al. [17] proposed the OCR-VQA task and created the OCR-VQA-200K dataset with 207,572 book cover images and over 1 million question-answer pairs. Biten et al. [5] addressed limitations in scene text understanding with the ST-VQA dataset, specifically designed for questions requiring knowledge of text present in images. Moving beyond natural scenes, Mathew et al. [18] introduced DocVQA, focusing on document images like invoices and contracts, with 50,000 question-answer pairs based on 12,767 document images. Finally, Tanaka et al. [19] advanced document-based VQA with VisualMRC, an abstractive task featuring over 30,000 question-answer pairs and 10,000 diverse document images, requiring models to generate natural language responses based on comprehensive document understanding.
@@ 2.1.2 Visual Question Answering Datasets in Vietnamese
Dataset Type Image Source Annotation Method Language ViVQA [11] Non text MS COCO Semi-Automation Vietnamese EJVVQA [20] Non text Internet sources Human annotation English, Japanese, Vietnamese OpenViVQA [12] Text & Non text Internet sources Human annotation Vietnamese ViOCRVQA [13] Text Internet sources - bookcover Semi-Automation Vietnamese ViTextVQA(ours) Text Internet sources, hand-craft Human annotation Vietnamese
In developing ViTextVQA, we draw upon both English resources and existing Vietnamese VQA datasets to better understand Vietnamese language structure and cultural context. The ViVQA dataset by Tran et al. [11] represents the first Vietnamese VQA dataset, containing 10,320 images and 15,000 question-answer pairs, though its semi-automatic creation method results in questions that don’t fully capture Vietnamese linguistic nuances and omits text elements in images. The EVJVQA dataset introduced by Nguyen et al. [20] offers a unique multilingual collection of human-annotated visual question-answer pairs using images from Vietnam, comprising approximately 33,790 question-answer pairs and over 5,000 images, with 11,689 pairs in Vietnamese and the remainder translated into English and Japanese. This dataset served as the primary resource for the Multilingual Visual Question Answering task at VLSP 2022. Addressing the limitations of ViVQA, Nguyen et al. [12] created OpenViVQA, the first large-scale Vietnamese VQA dataset with open-ended answers, featuring 11,199 images with 37,914 question-answer pairs in natural language forms rather than as a classification task. Finally, the ViOCRVQA dataset by Pham et al. [13] focuses specifically on OCR-VQA, providing Vietnam’s largest VQA dataset with nearly 30,000 images and over 120,000 question-answer pairs centered on scene text from Vietnamese book covers, created through semi-automatic labeling. Referencing these Vietnamese datasets helps us create questions and answers that accurately reflect Vietnam’s cultural context and characteristics while overcoming existing limitations.
The existing Vietnamese VQA datasets have several common limitations. ViVQA lacks natural language fluency in its questions and does not include text in images, while EVJVQA misses out on scene text, overlooking valuable OCR-related information. OpenViVQA, though designed for open-ended answers, faces challenges in classification tasks, and ViOCRVQA, despite being large and specialized, focuses primarily on text from book covers, lacking diversity in image types. These limitations highlight the need for more comprehensive, diverse, and culturally relevant datasets to improve the performance of VQA models in the Vietnamese context.
@@ 2.2.3 Vision Transformer Model - Transformer Language Model Based Method
3 Task and Dataset
First, we define text-based visual question answering (text-based VQA) task (see Section 3.1). Then, we introduce a rigorous process of data annotation, involving the participation of sixteen native Vietnamese speakers. The process of building the dataset is presented in Section 3.2. The prominent and intriguing features of the dataset, which we have identified through the labeling process, are detailed in Section 3.3. This Section includes an in-depth analysis of elements such as Question, Answer, and Object.
3.1 Task Definition
The Text-based Visual Question Answering (Text-based VQA) task is a subdomain of Visual Question Answering (VQA), which lies at the intersection of Computer Vision (CV) and Natural Language Processing (NLP). The primary goal is to develop a model capable of answering questions based on information from images. Specifically, given an input consisting of an image and a related question, the model must generate an accurate answer by leveraging the visual data present in the image.
In the case of Text-based VQA, the requirement is more specific: the image must contain text, and the question-answering process must rely heavily on the textual content embedded within the image. The Text-based VQA task is defined as following.
  • Input of this task includes an image and a question.
    • Image: An image that contains the necessary information for the model to analyze in order to answer the question. The image can belong to various categories, such as landscapes, people, or street scenes, but the crucial condition is the presence of text within the image.
    • Question: A question related to the content of the image. The question should be designed to align with the task’s purpose, which is to derive an answer grounded in the textual content present in the image.
  • Output of this task is a suitable answer to the question.
    • Answer: Given the input image and question, the model generates a corresponding answer by utilizing the textual information embedded in the image.
@@ 3.2.3 Question-Answer Creation
Next, we recruited 16 undergraduate students. Then, we trained them to understand the annotation guideline and asked them to annotate on about 100 test images. After ensuring that all students understood and applied the annotation guideline correctly, we appointed one person from our team for every four students to act as a supervisor. Annotation supervisors are responsible for monitoring and supporting the annotators, ensuring that everyone is performing according to the required process and quality.
@@ 3.3 Dataset Analysis
3.3.1 Question Length
@@
3.3.2 Answer Length